Operations Report – March 9th

Crew 206 Operations Report 09 – MAR – 2019

SOL: 13

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): 13,5 hours

Solar— SOC % – Turned off at 80%, Turned on at 73% (18h44)

Diesel Reading – 60%
Propane Reading – 76%

Ethanol Free Gasoline – 2 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 375 gallons

Water differential (static tank) – 13 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 226 gallons

Static to Loft Pump used – Yes

Water Meter: 0141711.2 gallons

Toilet tank emptied: Yes

Deimos rover used: No, still not functional
Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Breaking system tested.
Hours: 53,1 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

Curiosity rover used: Used
Hours: 78.2 hours

Beginning charge: 90%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Used
Hours: 72,2 hours

Beginning charge: 100%

Ending charge: 72 %

Currently charging: Yes

Notes on rovers: Opportunity’s brake fluid is empty, it looks like the pipeline has a leak. Atila is working on it; more information will follow.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Used by Staff.

Staff Crew Car used and why, where? Off-Campus

General notes and comments: Nothing to report.

Summary of internet: Internet access fixed! Thank you again, Bernard.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing more to report.

EVA Report – March 9th

Crew 206 EVA Report 09-MAR-2019
EVA #9

Author: Norbert Pouzin

Purpose of EVA:

The purpose of this EVA was to test our new ‘SUPAERO spacesuit’ on a short exploration EVA, at Kissing Camels Ridge West.

After our EVA #6 where we tested it near the Hab and made basic nominal checking, we wanted to go further.

What will be tested with the new spacesuit:
– Ability to drive a rover
– Ability to walk for more than an hour
– Testing its autonomy
– Making little sample returns (rocks), to test the storage of the spacesuit

Start time: 9:11 am

End time: 10:49 am

Narrative: This morning we decided that this EVA would be safe to conduct, as there were no wind nor rain. So we left airlock at 9:11 to go to Kissing Camels Ridges, with Curiosity and Spirit.

This EVA was made to test our new prototype, the ‘SUPAERO spacesuit’.

Thus, I wore it for the EVA, as I had already worn it for the last EVA with the spacesuit (EVA#6), to do several tests.

The first thing we tested was my ability to conduct a rover, here Spirit, with the spacesuit. All went well since we had added a little cushion to support the weight of my backpack.

Then, we arrived at 9:29 at Kissing Camel Ridge, and started exploration, for more than an hour. I was able to walk that long, supporting the backpack, and the prototype proved us that it had enough autonomy to do such exploration.

It appeared that with this spacesuit, there is no fog at all on the helmet, but also that it is more tiresome to wear.

Aurélien also did some samples returns, for an educational project for a primary French school.

We came back to the hab at 10:49, after a successful EVA.

Destination: Kissing Camel Ridge

Coordinates: 518300/4249300 (Kissing Camels Ridges)

EVA Participants: Norbert Pouzin (GHO, EVA leader), Aurélien Mure (XO), Jeremy Auclair (CMD), Cerise Cuny (HSO).

Roads and routes per MDRS Map: Entrance Road, Cow Dung Road

Mode of travel: Driving and Walking

Greenhab Report – March 9th

Crew 206 Green Hab Report – 09-MAR-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (3 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.6°C

High: 28.9°C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 6.5 gallons

Water in Blue Tank: 226.0 gallons

Crops watered at 0830 hours (2.2 gallons) and 1830 hours (4.3 gallons):


1) 08.30 AM 19.1° C / 28 %
2) 11.30 AM 22.5° C / 26 %
3) 15.00 PM 26.1° C / 20 %
4) 18.30 PM 21.6° C / 22 %

Changes to crops: Tomatoes planted last week are starting sprouting (3 seeds sprouted)!


Since we put the mercury thermometer in the Greenhab, it is easier to monitor the heat. I opened the door from 0830 to 1130, and it was not really hot in the afternoon.

I did my experiment this afternoon and made some interesting observations about my radishes, I will see in the next days how it goes.

Harvest: None.

Support/supplies needed: None.

Journalist Report – March 9th

Crew 206 – 03/09/2019

Sol 13

Authors : Jérémy Auclair, Commander and Benjamin Auzou, Journalist

"Martian Parliament"

Life is often made of compromises and debates. It’s important for the development of a society that people listen to each other’s points of view and take things into account to make the best choices. Even more in a mission like ours were a little number of people have to live together in a confined environment and take important decisions. It’s essential for everyone to share his/her experience and opinion to help the community.

The question of Mars and the exploration of outer space are also an open debate: How to deal with planets’ exploitation? Which laws apply to Mars? What authority will Earth have on a Mars colony? How to define the status of the settlers? There are many unanswered questions that have to be addressed before planning the first manned mission to Mars.

Since the beginning of the mission, we have spent more and more time debating, on various subjects and with different and interesting opinions. It’s a very entertaining activity, with an open mind it’s a great opportunity to learn new facts and to take a step back and see a little more of the whole picture. We have covered a wide range of
subjects. But one thing we have noticed since the middle of our second week is that our reactions have changed during our debates.

All crewmembers are getting more and more emotionally engaged in the debates, react stronger to the opposition and have more difficulty to overcome differences of opinions or ideology. We are all conscious of the effect confinement and isolation have on our social interactions, we see that we react differently to things we are very used to, but it is
still hard to control our responses. But we have prepared ourselves for that, we managed a wide range of situations before the mission that linked us strongy together, and we have the great tendency of laughing things out, which helps to take a step back.

Today was full of successes: we completed the second full Teleop
session (and first in the planned configuration), Norbert observed the
first results of his radiations and music on plants experiments,

Gaspard managed to debug the EVA tracking app he is working on, I baked two delicious cakes, we realized the first long EVA with our spacesuit and Cerise’s counting on Aquapad were interesting today.

However, a disastrous event broke that series of achievements: Gaspard tried to prepare a corn bread, and he is definitely better at engineering than baking…

Sol Summary – March 9th

Crew 206 March 9th, 2019
Sol 13

Summary Title: TELEOP catch-up

Author: Jeremy Auclair

Mission Status: All nominal

Sol Activity Summary: It is getting harder and harder for me to wake up in the morning, but I still stick to the morning routine, as all of the crew does. Today is the last working day of our week, and we worked well.

The EVA went beautifully. We explored the east and west Kissing Camels ridges and took great pictures. The EVA was also meant to continue the tests of our prototype simulation space suit. The little upgrades that were made improved the suit quite a bit: it was less hard to carry, ventilation and sun protection worked better and the battery lasted more than long enough for the whole EVA.

This afternoon also saw the completion of the TELEOP second week: we managed to gather additional data to catch up on the sessions that were missed, and I haven’t encountered any additional issues with the physical rover (great difference with last year’s sessions). We ended this work week with the VR session, that also allowed us to talk about the videos we had seen, and opinions differ regarding what video was preferred.

Look ahead plan: Tomorrow will be a rest day, we will still have to
work on our rotation video and on a few experiments, but it will be
calmer than today.

Anomalies in work : None

Weather : Cloudy all day with some sunshine.

Crew physical status : Fine

EVA : Kissing Camels

Reports to be filed : GreenHab, Journalist, Operation, EVA Report

Support Requested : None

Copyright © The Mars Society. All rights reserved. | Main Site